3. (30 points) In an RLC circuit with a resistance of 100 ohm connected to an AC power source with an amplitude of 80 V, the current reaches a maximum prior to the supplied voltage, which is only at 75% of its maximum when the current is at a maximum. (a) (11 points) What is the phase constant of the circuit? (b) (8 points) Is the circuit more inductive or more capacitive? (c) (11 points) What is the average power dissipated in this circuit?
Answer #1

current leads the voltage so this is capacitive circuit.

Xc > XL


(A) phase constant = 0.25 ( 90 deg ) = 22.5 deg  


(B) more capacitive


(C) cos(phi) = R / Z

Z = 100 / cos22.5 = 108.2 ohm

I = 80 / 108.2 = 0.739 A

Irms = I / sqrt(2) = 0.523 A


P_avg = Irms^2 R = 27.3 W